OpenStudy (anonymous):

A druggist has two mixtures of a certain chemical and water, one containing 10% of the chemical and the other containing 40% of the chemical. How much of each mixture should he use to make 10 ounces that contain 25% of the chemical? How many ounces of chemical are in the final mixture?

OpenStudy (anonymous):

If x is the amount of 10% mixture used, then which expression represents the ounces of chemical contained in it?

OpenStudy (anonymous):

the one that says \(.1\times x\)

OpenStudy (anonymous):

To the first question?

OpenStudy (anonymous):

Is f x is the amount of 10% mixture used, then which expression represents the ounces of chemical contained in it? = .1 X x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks. Do you know How many ounces of chemical are in the final mixture?

OpenStudy (anonymous):

or just \(.1x\)

OpenStudy (anonymous):

well if there are a total of 10 ounces, and \(x\) of them are 10% chemical, then the the rest is \(10-x\) and it is 40% chemical for a total of \(.1x+.4(10-x)\) ounces of chemical you want this to be 20% of 10 i.e. 2.5 ounces of chemical, so solve \[.1x+.4(10-x)=2.5\] for \(x\)

OpenStudy (anonymous):

sorry 25% of 10 is 2.5 i meant to say

OpenStudy (anonymous):

easier to solve if you multply both sides by ten and start with \[x+4(10-x)=25\]

OpenStudy (anonymous):

I get 5, but that's not an option in the answers

OpenStudy (anonymous):

Theres 1, 2.5, and 7.5

OpenStudy (anonymous):

\[x+40-4x=25\] \[-3x+40=25\] \[-3x=-15\] \[x=5\] is what i get too

OpenStudy (anonymous):

Damn

OpenStudy (anonymous):

ooooh hold on

OpenStudy (anonymous):

you solved the problem correctly, but they did not ask for a solution

OpenStudy (anonymous):

It asked How many ounces of chemical are in the final mixture?

OpenStudy (anonymous):

they must be trying to get you do do it step by step there are two questions in your post, but here is the one at the end "How many ounces of chemical are in the final mixture? "

OpenStudy (anonymous):

That's what I don't know how to solve

OpenStudy (anonymous):

that is easy final mixture is 25% of 10 is 2.5 ounces of chemical

OpenStudy (anonymous):

Thanks!! I got a 100 percent on that quiz because of you!

OpenStudy (anonymous):

not much to that problem, take 25% of ten

OpenStudy (anonymous):

yay